# RaSeRec: Retrieval-Augmented Sequential Recommendation

### v2

@@ Appendix H Study on User Interaction Frequency (RQ6)
Table 8: Performance comparison (@10), where the best results are boldfaced and the second-best ones are underlined.
(a) Results on Beauty w.r.t. HR@5 and NDCG@5.
(b) Results on Sports w.r.t. HR@5 and NDCG@5.
Figure 8: Model performance w.r.t. noise ratio on Beauty and Sports datasets. The bar represents HR@5 or NDCG@5, while the line represents the percentage of performance degradation compared to the test user sequences without added noise.
Appendix I Robustness against Noise Data (RQ7)
RQ7: Can RaSeRec perform robustly against the data noise issue?
In real-world scenarios, recommender systems usually suffer from the noisy interaction issue Gao et al. (2022); Tian et al. (2022), e.g., a user clicked an item by mistake. As such, it is necessary to verify RaSeRec’s robustness against noisy data. To simulate this scenario, we train models with full training data and randomly add a certain ratio (i.e., 10%, 20%, and 30%) of negative items into each test user sequence. The experimental results are shown in Figure 8, where models with retrieval augmentation mention “w/ aug” otherwise “w/o aug”.
From the experimental results, we observe that adding noises significantly degrades the performance of models with or without retrieval augmentation. However, ‘w/ aug’ consistently outperforms ‘w/o aug’ under any noise ratio in terms of HR@5 and NDCG@5. Moreover, the blue line is always below the green line, as shown in Figure 8, demonstrating that ‘w/ aug’ can endow the model backbone with more robustness against the noisy interactions during inference. For example, on the Beauty and Sports datasets, ‘w/ aug’ with a 20% noise ratio outperforms ‘w/o aug’ with a 10% noise ratio. Moreover, on the Sports dataset, ‘w/ aug’ drops 14.7% of its original performance, while ‘w/o aug’ drops 18.4%, in terms of NDCG@5, when both have a 10% noise ratio. This fully implies that with retrieval augmentation, RaSeRec can retrieve useful sequential patterns as references to alleviate the influence of noisy interactions and make a suitable recommendation.
In a nutshell, we suggest performing retrieval augmentation to alleviate the negative influence of noisy interactions.
Appendix J Learning Algorithm of RaSeRec
Algorithm 1 demonstrates the learning algorithm of the proposed RaSeRec framework. Generally, the learning process can be divided into three stages: (1) Initialization Stage (lines 1-4), which initializes the embedding table, the position encoding, and the model backbone; (2) Collaborative-based Pre-training Stage (lines 5-8), which learns to recommend items and retrieve memories based on collaborative signals; and (3) Retrieval-Augmented Fine-tuning Stage (lines 9-14), which learns to leverage retrieved memories to augment user representation for higher quality.
Algorithm 1 Learning Algorithm of RaSeRec
1:The behavior sequence of all users $\mathcal{S}=\{s_{u_{1}},s_{u_{2}},...,s_{u_{|\mathcal{U}|}}\}$, each user sequence defined as $s_{u}=\{v_{1}^{(u)},v_{2}^{(u)},...,v_{|s_{u}|}^{(u)}\}$; The maximum input sequence length $T$; The number of Transformer layers $L$; The number of retrieved memories $K$; The control coefficients $\alpha$ and $\beta$.
2:The pre-trained model backbone $\mathrm{SeqRec}(\cdot)$; The built memory bank $\mathcal{M}$; The fine-tuned RAM.
3:#1. Initialization Stage
4:Build the embedding table $\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times d}$ and an learnable position encoding matrix $\mathbf{P}\in\mathbb{R}^{T\times d}$;
5:Build an $L$-layer Transformer encoder with unidirectional attention $\mathrm{Trm}(\cdot)$;
6:Initialize the parameters of the model backbone $\mathrm{SeqRec}(\cdot)$, including $\mathbf{V}$, $\mathbf{P}$, and $\mathrm{Trm}(\cdot)$;
7:#2. Collaborative-based Pre-training Stage
8:Perform recommendation learning $\mathcal{L}_{rec}$ with Equ. (5) to endow the model backbone with the ability to generate the next item based on the input user sequence;
9:Construct positive and negative pairs for retrieval training with Equ. (6) and Equ. (7);
10:Perform retrieval training $\mathcal{L}_{ret}$ with Equ. (8) to endow the model backbone with the ability to retrieve memories with the same (or similar) preference as the input user sequence;
11:#3. Retrieval-Augmented Fine-tuning Stage
12:Build the memory bank $\mathcal{M}$, where each entry consists of a user representation and the corresponding target item embedding;
13:Build the retrieval-augmented module (RAM) and freeze the model backbone $\mathrm{SeqRec}(\cdot)$;
14:Retrieve top-$K$ memories (i.e., $\{{\mathbf{h}}_{k}\}_{k=1}^{K}$ and $\{{\mathbf{v}}_{k}\}_{k=1}^{K}$) from the memory bank $\mathcal{M}$ via the Faiss library;
15:Augment the user representation with the retrieved memories with Equ. (9), Equ. (10), and Equ. (11);
16:Perform retrieval-augmented fine-tuning $\mathcal{L}_{raft}$ with Equ. (12), making the model quicker to accommodate preference drifts and easier to recall long-tailed patterns;
17:return The pre-trained model backbone $\mathrm{SeqRec}(\cdot)$; The built memory bank $\mathcal{M}$; The fine-tuned RAM.

### v3

@@ Appendix C Baselines
Appendix D Metrics
[…]
    (14) $\begin{split}&\text{NDCG@}N=\frac{1}{|\mathcal{U}|}\sum_{u\in\mathcal{U}}\sum_{j=1}^{N}\\ &\delta(v_{t+1}=R(u)_{j})\frac{1}{\log_{2}(j+1)},\end{split}$
    where $R(u)_{j}$ denotes the $j$-th recommended item for user $u$.
Table 9. Hyperparameter settings, where * denotes different hyperparameter settings for different datasets.
Configuration Value
number of Transformer layers $L$ 2
number of heads 2
hidden size $d$ 64
maximum sequence length $L$ 50
number of clusters $k$ 128
number of clusters accessed 1
control coefficient $\alpha$* $[0.0,1.0]$
control coefficient $\beta$* $[0.0,1.0]$
number of retrieved memories $K$* $[5,55]$
learning rate 0.001
optimizer Adam
mini-batch size 1024
Appendix E Implementation Details
We use the Adam optimizer (Kingma and Ba, 2015) to optimize model parameters with the learning rate of 0.001, the mini-batch size of 1024, $\beta_{1}=0.9$, and $\beta_{2}=0.999$, where the maximum number of epochs is set to 100 for both pre-training and fine-tuning stages. We train models with the early stopping strategy that interrupts training if the HR@10 result on the validation set continues to drop until 10 epochs. For RAM, we tune $\alpha$, $\beta$ and $K$ within the ranges of $\{0.0,0.1,0.2,...,1.0\}$, $\{0.0,0.1,0.2,...,1.0\}$ and $\{5,10,15,...,55\}$, respectively, referring to Appendix F for more details. As for Faiss, we set the number of clusters $k$ to 128, set the number of clusters accessed as 1, and adopt the cosine similarity as the measurement to retrieve relevant memories. As for the sequence encoder, we adopt a two-layer Transformer and set the head number as 2 for each self-attention block, where the hidden size $d$ and the maximum sequence length $L$ are set to 64 and 50 respectively. Additionally, for Non-SeRec and Vanilla SeRec methods, we use results reported by DuoRec (Qiu et al., 2022) as these methods already have generally acknowledged experimental results on the three benchmark datasets used in this paper. We ran each experiment five times and reported the average. Table 9 summarizes the detailed hyper-parameter settings.
Figure 7. Parameter sensitivity w.r.t. the control coefficients $\alpha,\beta$ and the number of retrieved memories $K$, where we plot the error bar with the mean $\pm$ $\gamma\times$standard deviation and $\gamma$ is set as 500×500$\times$500 ×the scale interval of the y-axis.
(a) Results on the Beauty dataset w.r.t. HR@5 and NDCG@5.
(b) Results on the Sports dataset w.r.t. HR@5 and NDCG@5.
Figure 8. Performance comparison over different user groups between the model with retrieval augmentation and without retrieval augmentation. The bar represents HR@5 and NDCG@5, while the line denotes the performance improvement percentage of “w/ aug” compared to “w/o aug”.
Appendix F Parameter Sensitivity (RQ7)
Thirdly, we observe that the performance of RaSeRec reaches the peak to different $K$ across three datasets, e.g., 20 is the best value for the Beauty dataset. This demonstrates the effectiveness of augmenting user representation via retrieved explicit memories and manifests that setting a suitable number of retrieved memories can considerably enhance performance. The results on the three datasets follow similar trends, where the performance gradually improves as $K$ increases and then starts to degrade when further increasing $K$. This phenomenon is also observed in long-context RAG (Jin et al., 2024; Yue et al., 2024), where the generative modeling performance initially increases and then declines when consistently increasing the number of retrieved passages. The main reason is that increasing the number of retrieved memories will inevitably introduce irrelevant memories (i.e., “noise”) that can mislead the retrieval-augmented module. In conclusion, we suggest tuning $K$ in the range of $10\sim 30$ for the Beauty and Sports datasets, and $30\sim 50$ for the Clothing dataset.
Appendix G Study on User Interaction Frequency (RQ8)